### v1

# ParsVoice: A Large-Scale Multi-Speaker Persian Speech Corpus for Text-to-Speech Synthesis

Abstract
Persian Language, despite being spoken by over 100 million people worldwide, remains severely underrepresented in high-quality speech corpora, particularly for text-to-speech (TTS) synthesis applications. Existing Persian speech datasets are typically smaller than their English counterparts, which creates a key limitation for developing Persian speech technologies. We address this gap by introducing ParsVoice, the largest Persian speech corpus designed specifically for TTS applications. We created an automated pipeline that transforms raw audiobook content into TTS-ready data, incorporating components such as a BERT-based sentence completion detector, a binary search boundary optimization method for precise audio-text alignment, and multi-dimensional quality assessment frameworks tailored to Persian. The pipeline processes 2,000 audiobooks, yielding 3,526 hours of clean speech, which was further filtered into a 1,804-hour high-quality subset suitable for TTS, featuring more than 470 speakers. ParsVoice is the largest high-quality Persian speech dataset, offering speaker diversity and audio quality comparable to major English corpora. The complete dataset has been made publicly available to accelerate the development of Persian speech technologies and to serve as a template for other low-resource languages. The ParsVoice dataset is publicly available at ParsVoice.
Index Terms—  Text-to-Speech, Persian Speech Corpus, Low-resource Languages, Multi-speaker Dataset, Speech synthesis
1 Introduction
Deep learning has fundamentally transformed artificial intelligence, yet it has also introduced models that are increasingly data-hungry. As transformer architectures [1] and generative models, rapidly advance, the scarcity of high-quality training data has become even more pronounced, especially for low-resource languages. Persian, although spoken by more than 100 million people worldwide, remains significantly underrepresented in speech and text corpora when compared to high-resource languages such as English.
Text-to-speech (TTS) systems present unique data requirements that differ substantially from automatic speech recognition (ASR) systems. While ASR models can benefit from training on noisy, real-world data that reflects actual usage conditions, TTS models require exceptionally clean and precisely aligned audio-text pairs to generate natural-sounding speech. This requirement makes high-quality TTS datasets significantly more challenging and expensive to create.
The need for high quality data is particularly evident in low-resource languages, where models tend to underperform such as Persian. The shortage of Persian data is not limited to speech but extends to text and images, creating cascading effects across multiple areas of Persian language processing. Speech-to-text alignment systems, optical character recognition (OCR) models, and other deep learning applications in Persian all suffer from insufficient training resources. This scarcity has slowed the development of robust Persian language technologies and further widened the digital divide between Persian and well-resourced languages.
[…]
Fig. 1: Overview of the proposed pipeline.
2 Related Work
The development of speech datasets has been predominantly driven by English resources, with corpora such as LibriSpeech [2], LJSpeech [4], and VCTK [5] establishing long-standing benchmarks for ASR and TTS research. Multilingual datasets, including Multilingual LibriSpeech [6] and VoxPopuli [8], have expanded coverage to more than twenty languages, but remain largely skewed toward European languages. Similarly, Common Voice [3] offers crowdsourced data for over 100 languages, though quality and consistency vary considerably across linguistic communities.
Among languages geographically and linguistically related to Persian, resource availability is uneven. Kurdish has benefited from efforts such as the KASET corpus [9], while Pashto and Dari contributions are limited to small-scale subsets of Common Voice. In contrast, neighboring languages such as Turkish [10] and Arabic [12] enjoy more comprehensive and increasingly standardized datasets.
However, Persian speech processing, continues to face severe resource scarcity. The Persian portion of Common Voice provides only ASR-suitable data, lacking the precise alignment and quality required for TTS. Several other Persian datasets exist but suffer from significant limitations: DeepMine+ [27] offers over 480 hours of ASR data from 1850+ speakers but is commercially restricted under proprietary licensing. The FarsSpon dataset [31] provides 530+ hours from over 5300 speakers but remains commercially unavailable. For TTS specifically, ArmanTTS [28] contains approximately 9 hours from a single speaker but has not been made publicly available, ManaTTS [18] was recently introduced as a publicly accessible single-speaker Persian corpus, comprising approximately 86 hours of audio; however, it is limited to a single speaker. Despite these efforts, the Persian speech processing field still lacks the comprehensive, high-quality datasets available for higher-resource languages.
The fundamental limitation lies in scale: existing Persian TTS corpora are much smaller than their English counterparts, and most are restricted to single-speaker settings. This scarcity significantly constrains the development of robust, multi-speaker TTS systems, underscoring the need for scalable, automated approaches to corpus construction that can meet the demands of the 100+ million Persian speakers worldwide.
3 Automated Corpus Construction Pipeline
We introduce a pipeline that transforms raw audiobook recordings into a structured, high-quality speech and text corpus through interconnected stages designed to maximize data quality. The pipeline addresses key challenges in Persian speech data creation: maintaining sentence integrity, ensuring audio-text alignment accuracy, and scaling to thousands of hours of content.
3.1 Data Collection and Source Selection
We selected IranSeda (book.iranseda.ir) as our primary data source based on several critical considerations:
Content Diversity: The platform hosts over 3,800 audiobooks across many categories including classical literature, contemporary fiction, science, philosophy, history, and educational content, ensuring broad lexical and stylistic coverage essential for robust TTS training.
Professional Quality: Content is produced by narrators in controlled recording environments at 44.1 kHz sampling rate, providing consistent audio quality crucial for neural TTS model training.
Open Availability: All audiobook materials from IranSeda are freely accessible to the public without copyright restrictions, allowing the resulting dataset to be distributed, downloaded, and used for both academic research and practical speech technology development.
3.2 Intelligent Audio Segmentation
Raw audiobook files typically span several hours each, require segmentation into appropriate chunks, while preserving sentence integrity—a non-trivial requirement that existing alignment tools fail to handle effectively, particularly for Persian language processing with imperfect audio-text alignment scenarios.
To ensure dataset quality through the preservation of complete sentences, we developed a sentence completion detection model by fine-tuning ParsBERT [7], a transformer-based model for the Persian language. This model was trained on a custom-built dataset of complete and incomplete sentences to automatically identify and filter sentence fragments, achieving an F1 score of 97.4. The classifier is integrated into a three-phase segmentation pipeline designed to maintain linguistic coherence.
Phase 1: Acoustic Boundary Detection. We first apply WebRTC [34] Voice Activity Detection (VAD) with aggressiveness level 1, which achieved the highest sentence completion rate in our empirical evaluation (61.4% compared to 49.5% for Whisper). This step detects candidate speech segments by locating silence-based boundaries in the audio signal.
Phase 2: Transcription and Alignment. Each candidate segment is transcribed using the free Google Speech-to-Text API, chosen for its lower word error rate (WER) on Persian speech. On the Persian Common Voice dataset—which is notably noisy and challenging—the API achieves a WER of 25.2%, compared to 31.75% reported for Whisper [35]. The resulting transcriptions provide a reliable basis for subsequent linguistic validation.
Phase 3: Linguistic Validation. Finally, each transcription is analyzed using our BERT-based sentence completion classifier. Segments identified as incomplete undergo iterative boundary extensions in 0.1-second increments (up to 5 seconds), looping between ParsBERT and the ASR model until the transcription meets the completeness criteria or is rejected due to length constraints.
3.3 Boundary Optimization Algorithm
Even with accurate transcription, audio segments may contain unwanted silence, background noise, or acoustic artifacts at boundaries that degrade TTS model performance. Our boundary optimization algorithm employs binary search to determine optimal trimming points for both segment start and end boundaries.
The boundary optimization follows a two-stage search strategy combining binary refinement with linear fine-tuning:
1. Initial Adjustment: We begin by trimming 3 seconds from both the start and end of the segment, followed by re-transcription.
2. Transcription Stability Check: If the new transcription differs from the original, the removal is considered excessive, and the boundary is readjusted.
3. Binary Search Refinement: The algorithm then applies a binary search procedure, iteratively halving the adjustment interval to rapidly narrow the boundary location.
4. Linear Precision Search: Once the optimal boundary is found using binary search, we switch to a linear search with 0.1-second steps for ranges that cannot be resolved by binary search, ensuring exact alignment.
This approach ensures that each segment contains only the essential speech content while maintaining transcription accuracy.
3.4 Multi-Dimensional Quality Assessment
Quality control is critical for TTS training data it also can help for ASR training, as low-quality samples can degrade model performance. We implement comprehensive assessment across audio and text dimensions.
3.4.1 Persian Text Quality Framework
The Persian Text Quality Framework evaluates transcriptions to ensure they are well-formed, representative of Persian, and suitable for TTS training. Each sentence is assessed across multiple dimensions, which are combined into a weighted total score. The main criteria include character quality, which measures the proportion of valid Persian characters and digits while penalizing excessive use of foreign characters; length quality, which evaluates whether the sentence length in words and characters falls within an ideal range, with very short or very long sentences reducing the score; repetition score, which penalizes excessive word repetition and rewards higher lexical diversity; and phonetic coverage, which favors sentences that encompass a broad range of Persian characters and phonemes, including both vowels and consonants, to maximize phonetic diversity.
Each metric is first normalized to the range $[0,1]$ and then aggregated using predefined weights, empirically determined, to compute the total score. Sentences are categorized based on this score: those with a score of 0.7 or higher are considered high quality and recommended for TTS training; scores between 0.5 and 0.7 indicate mid-quality sentences that are acceptable but may require review; and sentences scoring below 0.5 are deemed low quality and not recommended for use.
3.4.2 Audio Quality Metrics
The audio quality framework assesses recordings for overall clarity, absence of distortions, and suitability for speech processing tasks. Each file is automatically analyzed, and a composite quality score is calculated based on standard audio metrics, including signal-to-noise ratio, dynamic range, spectral features, MFCC variance, clipping, silence, background music presence (using inaSpeechSegmenter [32]), and duration. Each metric is first normalized to the range $[0,1]$ and then aggregated using empirically determined weights to produce a final score scaled from 0 to 1. Recordings with scores above 0.9 are considered high quality, scores between 0.75 and 0.9 indicate acceptable quality, and scores below 0.75 are deemed low quality. Higher scores reflect cleaner, more usable recordings, while lower scores correspond to files affected by noise, distortion, silence, or background music.
3.5 Speaker Identification
Our metadata contained information about the main narrators of each book; however, many entries (approximately 40% of the dataset) lacked narrator names. Additionally, some books had multiple narrators without a clear distinction. Consequently, the exact number of speakers in the audio was unknown, making precise speaker identification necessary. To assign consistent speaker labels, we used a two-stage identification pipeline based on ECAPA-TDNN embeddings [33].
3.5.1 Local Speaker Diarization
For each audiobook, we perform an independent clustering process. The embeddings undergo preprocessing, including outlier removal, L2 normalization, and dimensionality reduction via UMAP. To determine the number of speakers ($k^{*}$), we employ a robust ensemble estimation method. We calculate the optimal $k$ suggested by Silhouette, Calinski-Harabasz, and Davies-Bouldin scores, as well as an estimate from HDBSCAN. The final $k^{*}$ is chosen by a consensus vote among these methods, falling back to the median if no consensus is reached.
An ensemble of clustering algorithms (e.g., Agglomerative, Spectral) is then applied, and the best result is selected based on the highest silhouette score. Each segment is assigned a confidence score $c_{i}$, which is a weighted combination of its individual silhouette score ($s_{i}$) and its relative distance to its assigned cluster centroid versus the nearest alternative. Segments with low confidence are filtered out.
3.5.2 Global Speaker Identification
To identify speakers across the entire corpus, we first represent each local speaker cluster by its medoid embedding, chosen for its robustness to outliers. A pairwise cosine similarity matrix is computed between all local speakers from all books. We convert this to a distance matrix (distance = 1 - similarity) and apply HDBSCAN to group local speakers into globally consistent identities, filtering clusters below a minimum confidence threshold. These global speaker IDs are then used to label the final dataset.
3.6 Final Data Cleaning and Preparation
To prepare the final dataset suitable for TTS model training, we created a high-quality subset specifically for TTS applications. In this subset, we removed audio files with quality scores below 0.8 and text segments with scores below 0.5. Additionally, we fine-tuned a ParsBERT model on a custom-built corpus for Persian punctuation restoration, achieving an F1 score of 91.33%. This model was subsequently applied to our dataset to reconstruct missing punctuation, ensuring all text segments are properly punctuated and structurally complete.
Table 1: ParsVoice Dataset Statistics
Metric Before Filtering After TTS Filtering
Total Hours 3,526.4 1,803.9
Segments 2,603,045 1,147,718
4 ParsVoice Corpus Analysis
4.1 Collection and Processing Results
Out of 3,807 books (9,538 hours), we fully processed 2,000. Our automated pipeline generated 5,158,344 initial audio segments. After removing empty segments, 3,321,212 segments remained.
Table 2: Selected Persian Text-to-Speech Datasets
Dataset Size (Hours) Speakers
ParsVoice (Ours) 1,804 470+
ManaTTS [18] 86 1
DeepMine Multi-TTS [19] 120 67
ArmanTTS [28] 9 1
AmerAndish [20] 21 1
ParsiGoo [25] 5 6
Persian Speech [26] 2.5 1
Quality Assessment. The boundary optimization algorithm removed 442.73 hours (11.2%) of unwanted silence and artifacts. Overall, 81.0% of segments required trimming at the start, while 50.4% required trimming at the end. The average segment duration is 5.49 seconds, which is optimal for TTS training.
Linguistic Analysis. Text analysis reveals substantial linguistic diversity, with 267,965 unique words from 25,499,474 tokens. The dataset has an average of 9.8 words per sentence and 3.8 characters per word.
Speaker Statistics. In our metadata, some narrator names were missing, and some books had multiple narrators. Gender analysis based on the 470 available narrator names revealed that ParsVoice consists of approximately 33% female and 67% male narrators. Our speaker identification pipeline detected over 1,815 unique speaker instances across the entire dataset. When analyzing only the subset of audiobooks that have narrator metadata (470 unique narrator names), we found that our global speaker IDs achieve 97.0% consistency with the known narrator labels, indicating highly reliable speaker identification.
5 Conclusion
In this work, we address the scarcity of high-quality Persian speech datasets by introducing ParsVoice, the largest publicly available Persian dataset to date. ParsVoice consists of 1,804 hours of clean, segmented speech from 470+ distinct speakers suitable for TTS training, and an additional 2,000 hours of high-quality speech that can be used in a wide range of speech research applications.
Alongside the dataset, we provide a fully scalable and automated pipeline for dataset creation. This pipeline incorporates several key innovations, including a BERT-based model to ensure sentence completeness, a binary search algorithm for precise audio-text boundary optimization, and comprehensive multi-dimensional quality assessment frameworks specifically designed for Persian. Together, the dataset and the pipeline provide a valuable resource for advancing Persian speech research and TTS development.