### v2

# ParsVoice: A Large-Scale Multi-Speaker Persian Speech Corpus for Text-to-Speech Synthesis

Abstract
Existing Persian speech datasets are typically smaller than their English counterparts, which creates a key limitation for developing Persian speech technologies. We address this gap by introducing ParsVoice, the largest Persian speech corpus designed specifically for text-to-speech(TTS) applications. We created an automated pipeline that transforms raw audiobook content into TTS-ready data, incorporating components such as a BERT-based sentence completion detector, a binary search boundary optimization method for precise audio-text alignment, and audio-text quality assessment frameworks tailored to Persian. The pipeline processes 2,000 audiobooks, yielding 3,526 hours of clean speech, which was further filtered into a 1,804-hour high-quality subset suitable for TTS, featuring more than 470 speakers. To validate the dataset, we fine-tuned XTTS for Persian, achieving a naturalness Mean Opinion Score (MOS) of 3.6/5 and a Speaker Similarity Mean Opinion Score (SMOS) of 4.0/5 demonstrating ParsVoice’s effectiveness for training multi-speaker TTS systems. ParsVoice is the largest high-quality Persian speech dataset, offering speaker diversity and audio quality comparable to major English corpora. The complete dataset has been made publicly available to accelerate the development of Persian speech technologies. The ParsVoice dataset is publicly available at ParsVoice.
Index Terms—  Text-to-Speech, Persian Speech Corpus, Low-resource Languages, Multi-speaker Dataset, Speech synthesis
1 Introduction
As transformer architectures [1] and generative models, rapidly advance, the scarcity of high-quality training data has become even more pronounced, especially for low-resource languages. Persian, although spoken by more than 100 million people worldwide, remains significantly underrepresented in speech corpora compared to high-resource languages such as English.
Text-to-speech (TTS) systems present unique data requirements that differ substantially from automatic speech recognition (ASR) systems. While ASR models can benefit from training on noisy, real-world data that reflects actual usage conditions, TTS models require clean and precisely aligned audio-text pairs to generate natural-sounding speech. This requirement makes high-quality TTS datasets more challenging and expensive to create.
The need for high quality data is particularly evident in low-resource languages, where models tend to underperform such as Persian. The shortage of Persian data is not limited to speech but extends to text, creating cascading effects across multiple areas of Persian language processing. Speech-to-text alignment systems, optical character recognition (OCR) models, and other deep learning applications in Persian all suffer from insufficient training resources. This scarcity has slowed the development of robust Persian language technologies and further widened the digital divide between Persian and well-resourced languages.
We address the challenge of data scarcity in Persian speech processing by introducing ParsVoice, the largest and most comprehensive Persian speech corpus designed specifically for modern TTS applications. We develop a scalable, automated pipeline for transforming raw audiobook content into TTS-ready data, incorporating novel techniques for sentence-aware segmentation, boundary optimization, and Persian-specific quality assessment. The resulting corpus comprises 3,526 hours of speech from 470+ unique speakers, representing a 10× increase over previous Persian speech resources and achieving speaker diversity comparable to major English corpora. To validate the usability of ParsVoice, we fine-tuned XTTS—a zero-shot multilingual TTS model that operates directly on text without requiring phoneme representations—on our corpus, in contrast to traditional Persian TTS systems that rely on explicit phonetic transcription. Our model achieved a mean opinion score (MOS) of 3.6/5 and a Speaker Similarity Mean Opinion Score (SMOS) of 4.0/5, demonstrating that ParsVoice enables high-quality, phoneme-free Persian speech synthesis and opens new possibilities for rapid speaker adaptation in Persian. An overview of the complete pipeline is illustrated in Figure 1.
Fig. 1: Overview of the proposed pipeline.
2 Related Work
Speech dataset development has been dominated by English resources such as LibriSpeech [2], LJSpeech [3], and VCTK [4]. While multilingual efforts like Common Voice [5], Multilingual LibriSpeech [6], and VoxPopuli [7] have expanded to 20+ languages, they remain skewed toward European languages with variable quality across linguistic communities. In contrast, many widely spoken non-European languages—such as Persian—have received comparatively little attention, leaving researchers with limited and often inaccessible data. Common Voice’s Persian portion lacks the quality required for TTS. Existing Persian datasets suffer from critical limitations: DeepMine+ [8] provides 480+ hours from 1850+ speakers but is commercially restricted; for TTS specifically, DeepMine-Multi-TTS provides 120 hours across 67 speakers, ArmanTTS [9] contains  9 hours from a single speaker, and ManaTTS [10] offers 86 hours from one speaker. Recent Persian TTS training efforts include ManaTTS, achieving 3.74 and 3.9 MOS; DeepMine-Multi-TTS, achieving 3.94 and 4.12 MOS with two multi-speaker synthesizers; A variant of FastSpeech2 [11] trained on the DeepMine dataset achieved 3.95 MOS and 3.32 SMOS. Notably, some traditional approaches require explicit phoneme representations, adding complexity to the pipeline and limiting accessibility.
Nevertheless, the fundamental limitation remains the scale and accessibility: existing Persian TTS corpora are orders of magnitude smaller than English counterparts, mostly single-speaker, and often proprietary. This severely constrains multi-speaker TTS development, necessitating scalable, open-source corpus construction approaches.
3 Automated Corpus Construction Pipeline
We introduce a pipeline that transforms raw audiobook recordings into a structured, high-quality speech and text corpus through interconnected stages designed to maximize data quality. The pipeline addresses key challenges in Persian speech data creation: maintaining sentence integrity, ensuring audio-text alignment accuracy, and scaling to thousands of hours of content.
3.1 Data Collection and Source Selection
We selected IranSeda (book.iranseda.ir) as our primary data source based on several critical considerations. The platform hosts over 3,800 audiobooks across many categories, ensuring broad lexical and stylistic coverage essential for robust TTS training. Content is produced by professional narrators in controlled recording environments at 44.1 kHz sampling rate, providing consistent audio quality crucial for neural TTS model training. Importantly, all audiobook materials from IranSeda are freely accessible to the public without copyright restrictions, allowing the resulting dataset to be distributed, downloaded, and used for both academic research and practical speech technology development.
3.2 Intelligent Audio Segmentation
Raw audiobook files typically span several hours each, require segmentation into appropriate chunks, while preserving sentence integrity—a non-trivial requirement that existing alignment tools fail to handle effectively, particularly for Persian language processing with imperfect audio-text alignment scenarios.
To ensure dataset quality through the preservation of complete sentences, we developed a sentence completion detection model by fine-tuning ParsBERT [13], a transformer-based model for the Persian language. This model was trained on a custom-built dataset derived from [14], which contains complete Persian sentences. Incomplete sentences were synthetically generated by randomly removing up to the last five characters or words from the complete ones. identify and filter sentence fragments, achieving an F1 score of 97.4. The classifier is integrated into a three-phase segmentation pipeline designed to maintain linguistic coherence.
Phase 1: Acoustic Boundary Detection. We first apply WebRTC [15] Voice Activity Detection (VAD) with aggressiveness level 1. This step detects candidate speech segments by locating silence-based boundaries in the audio signal.
Phase 2: Transcription and Alignment. Each candidate segment is transcribed using the free Google Speech-to-Text API, chosen for its lower word error rate (WER) on Persian.
Phase 3: Linguistic Validation. Finally, each transcription is analyzed using our BERT-based sentence completion classifier. Segments identified as incomplete undergo iterative boundary extensions in 0.1-second increments (up to 5 seconds), looping between ParsBERT and the ASR model until the transcription meets the completeness criteria or is rejected due to length constraints.
3.3 Boundary Optimization Algorithm
Even with accurate transcription, audio segments may contain unwanted silence, background noise, or acoustic artifacts at boundaries that degrade TTS model performance. Our boundary optimization algorithm employs binary search to determine optimal trimming points for both segment start and end boundaries.
The boundary optimization follows a two-stage search strategy combining binary refinement with linear fine-tuning:
1. Initial Adjustment: We start by removing 3 seconds from both the beginning and end of each segment, then perform re-transcription.
2. Stability Verification: If the new transcription differs significantly from the original, the trimming is deemed excessive and must be reduced.
3. Binary Search Optimization: The algorithm employs binary search, iteratively halving the trimming interval to efficiently converge on the optimal boundary position. Binary search continues until re-transcription differs from the original.
4. Fine-Grained Linear Search: After binary search converges, a linear search with 0.1-second increments is applied to achieve precise boundary alignment where further binary search refinement is no longer effective.
This approach ensures that each segment contains only the essential speech content while maintaining transcription accuracy.
3.4 Text-Audio Quality Assessment
High quality speech is critical for TTS training data it also can help for ASR training, as low-quality samples can degrade model performance. We implement comprehensive assessment across audio and text dimensions.
3.4.1 Persian Text Quality Metrics
The Persian Text Quality Framework evaluates transcriptions to ensure they are well-formed, representative of Persian, and suitable for TTS training. Each sentence is assessed across multiple dimensions, which are combined into a weighted total score. The main criteria include character quality, which measures the proportion of valid Persian characters and digits while penalizing excessive use of foreign characters; length quality, which evaluates whether the sentence length in words and characters falls within an ideal range, with very short or very long sentences reducing the score; repetition score, which penalizes excessive word repetition and rewards higher lexical diversity; and phonetic coverage, which favors sentences that encompass a broad range of Persian characters and phonemes, including both vowels and consonants, to maximize phonetic diversity.
Each metric is normalized to $[0,1]$ and aggregated using empirical weights to compute a total score. Sentences are categorized as: high quality ($\geq 0.7$, recommended for TTS training), mid-quality (0.50.5-$0.7$, acceptable but may need review), or low quality (<0.5<0.5, not recommended).
3.4.2 Audio Quality Metrics
The audio quality framework assesses recordings for overall clarity, absence of distortions, and suitability for speech processing tasks. Each file is automatically analyzed, and a composite quality score is calculated based on standard audio metrics, including estimated signal-to-noise ratio, dynamic range, spectral features, MFCC variance, clipping, silence, background music presence (using inaSpeechSegmenter [16]), and duration. Metrics are normalized to $[0,1]$ and empirically weighted to produce a final score. Recordings scoring above 0.9 are high quality, 0.75-0.9 are acceptable, and below 0.75 are low quality. Higher scores reflect cleaner recordings, while lower scores indicate noise, distortion, silence, or background music.
3.5 Speaker Identification
Our metadata contained information about the main narrators of each book; however, many entries (approximately 40% of the dataset) lacked narrator names. Additionally, some books had multiple narrators without a clear distinction. Therefore, the exact number of speakers in the audio was unknown, making precise speaker identification necessary. To assign consistent speaker labels, we used a two-stage identification pipeline based on ECAPA-TDNN embeddings [17].
3.5.1 Local Speaker Diarization
For each audiobook, embeddings are preprocessed (outlier removal, L2 normalization, UMAP reduction). The number of speakers ($k^{*}$) is determined by consensus vote among Silhouette, Calinski-Harabasz, Davies-Bouldin scores, and HDBSCAN estimates, defaulting to the median if needed.
Clustering algorithms (Agglomerative, Spectral) are applied, selecting the best result by silhouette score. Each segment receives a confidence score $c_{i}$ combining its silhouette score ($s_{i}$) and relative distance to its cluster centroid. Low-confidence segments are filtered out.
3.5.2 Global Speaker Identification
Each local speaker cluster is represented by its weighted centroid embedding. Pairwise cosine similarities between all local speakers are computed and converted to distances (1 - similarity). Agglomerative clustering with average linkage groups them into global identities, filtering low-confidence clusters. These global IDs label the final dataset.
3.6 Final Data Cleaning and Preparation
To prepare the final dataset suitable for TTS model training, we created a high-quality subset specifically for TTS applications. In this subset, we removed audio files with quality scores below 0.8 and text segments with scores below 0.5. Additionally, we fine-tuned a ParsBERT model on a custom-built corpus for Persian punctuation restoration[14], achieving an F1 score of 91.33%. This model was subsequently applied to our dataset to reconstruct missing punctuation, ensuring all text segments are properly punctuated and structurally complete.
Table 1: ParsVoice Dataset Statistics
Metric Before Filtering After TTS Filtering
Total Hours 3,526.4 1,803.9
Segments 2,603,045 1,147,718
4 ParsVoice Corpus Analysis
4.1 Collection and Processing Results
Out of 3,807 books (9,538 hours), we fully processed 2,000. Our automated pipeline generated 5,158,344 initial audio segments. After removing empty segments, 3,321,212 segments remained.
Table 2: ParsVoice Dataset in Comparison to Other Datasets
Dataset Size (Hours) Speakers
ParsVoice (Ours) 1,804 470+
ManaTTS [10] 86 1
DeepMine Multi-TTS [12] 120 67
Quality Assessment. The boundary optimization algorithm removed 442.73 hours (11.2%) of unwanted silence and artifacts. Overall, 81.0% of segments required trimming at the start, while 50.4% required trimming at the end. The average segment duration is 5.49 seconds, which is optimal for TTS training.
Linguistic Analysis. Text analysis reveals substantial linguistic diversity, with 267,965 unique words from 25,499,474 tokens. The dataset has an average of 9.8 words per sentence and 3.8 characters per word.
Speaker Statistics. In our metadata, some narrator names were missing, and some books had multiple narrators. Gender analysis based on the available narrator names revealed that ParsVoice consists of approximately 33% female and 67% male narrators. Our speaker identification pipeline detected over 1,815 unique speaker instances across the entire dataset. When analyzing only the subset of audiobooks that have narrator metadata, we found that our global speaker IDs achieve 97.0% consistency with the known narrator labels, indicating highly reliable speaker identification.
5 Evaluation: TTS Model Training
To validate ParsVoice for TTS applications, we fine-tuned XTTS [18], a state-of-the-art multi-lingual TTS model with zero-shot capabilities.
Training: A BPE model was trained, and 2,500 new Persian tokens were extracted from Copera and added to the GPT model vocabulary. The model was then fine-tuned with a batch size of 16 for 170,000 steps on the ParsVoice dataset.
Evaluation Protocol: We synthesized 90 samples using random sentences from unseen Persian speakers. Subjective evaluation by 40 raters assessed naturalness, speaker similarity, and text accuracy on a 1–5 scale. Objective metrics included WER/CER (using Google STT) and speaker similarity measured as the cosine similarity between reference and synthesized speech embeddings extracted using ECAPA-TDNN. Results: Table 3 demonstrates competitive performance with existing Persian zero-shot TTS systems. Our system achieved naturalness MOS of 3.6/5 and speaker similarity MOS of 4.0/5. Objective metrics showed WER of 22.57% and CER of 12.78%. However, these automatic metrics may underestimate quality due to limited Persian synthetic voice data in commercial STT systems. Human evaluation assessed how accurately the synthesized speech matched the input text, yielding 4.0/5 and confirming high quality. Additionally, speaker similarity of 80% based on ECAPA-TDNN embeddings demonstrates effective zero-shot capability.
Table 3: Comparison on Unseen Speakers
System MOS SMOS
XTTS + ParsVoice (Ours) 3.60 4.00
FastSpeech2 End-to-End [12] 3.72 4.02
FastSpeech2 Cascaded [12] 3.34 3.81
6 Conclusion
In this work, we address the scarcity of high-quality Persian speech datasets by introducing ParsVoice, the largest publicly available Persian dataset to date. ParsVoice consists of 1,804 hours of clean, segmented speech from 470+ distinct speakers suitable for TTS training, and an additional 2,000 hours of high-quality speech that can be used in a wide range of speech research applications.
Alongside the dataset, we provide a fully scalable and automated pipeline for dataset creation. This pipeline incorporates several key parts, including a BERT-based model to ensure sentence completeness, a binary search algorithm for precise audio-text boundary optimization, and comprehensive audio-text quality assessment frameworks specifically designed for Persian. Together, the dataset and the pipeline provide a valuable resource for advancing Persian speech research and TTS development.
We validated ParsVoice by fine-tuning XTTS, achieving competitive performance with naturalness and speaker similarity MOS scores of 3.6/5 and 4.0/5, respectively. These results confirm the corpus’s quality and its suitability for developing robust multi-speaker TTS systems, addressing a critical resource gap in Persian language technology.